Solve each system using substitution. Check your solution.
y=4x-8
y=2x+10

Answer by elima(1433)

y=4x-8
y=2x+10
Substitue one of the answers you have for y in the other equation;
(4x-8)=2x+10
4x-8-2x=10
2x=10+8
2x=18
x=9
Now plug that into the equation to get the answer for y;
y=4(9)-8
y=36-8
y=28
x=9,y=28
=)